### v1

# Leveraging Synthetic Data for Question Answering with Multilingual LLMs in the Agricultural Domain

Abstract
Enabling farmers to access accurate agriculture-related information in their native languages in a timely manner is crucial for the success of the agriculture field. Although large language models (LLMs) can be used to implement Question Answering (QA) systems, simply using publicly available general-purpose LLMs in agriculture typically offer generic advisories, lacking precision in local and multilingual contexts due to insufficient domain-specific training and scarcity of high-quality, region-specific datasets. Our study addresses these limitations by generating multilingual synthetic agricultural datasets (English, Hindi, Punjabi) from agriculture-specific documents and fine-tuning language-specific LLMs. Our evaluation on curated multilingual datasets demonstrates significant improvements in factual accuracy, relevance, and agricultural consensus for the fine-tuned models compared to their baseline counterparts. These results highlight the efficacy of synthetic data-driven, language-specific fine-tuning as an effective strategy to improve the performance of LLMs in agriculture, especially in multilingual and low-resource settings. By enabling more accurate and localized agricultural advisory services, this study provides a meaningful step toward bridging the knowledge gap in AI-driven agricultural solutions for diverse linguistic communities.
Leveraging Synthetic Data for Question Answering with Multilingual LLMs in the Agricultural Domain
Rishemjit Kaur1,2,*, Arshdeep Singh Bhankhar1,*, Surangika Ranathunga3,*, Jashanpreet Singh Salh1, Sudhir Rajput1, Vidhi1, Kashish Mahendra1, Bhavika Berwal1, Ritesh Kumar 1,2, 1CSIR-Central Scientific Instruments Organisation, India 2Academy of Scientific and Innovative Research (AcSIR), Ghaziabad, U.P., India 3School of Mathematical and Computational Sciences, Massey University, New Zealand * These authors contributed equally Correspondence: rishemjit.kaur@csio.res.in, s.ranathunga@massey.ac.nz
1 Introduction
Agriculture forms the backbone of many economies, specially for countries such as India, with agriculture being the primary occupation for a major part of the population (Ministry of Agriculture & Farmers Welfare, Government of India, 2024). Despite substantial advancements of Artificial Intelligence (AI) across healthcare, finance, and education (Li et al., 2024; Zhu et al., 2024), the agriculture sector has seen relatively limited integration of cutting-edge technologies such as Large Language Models (LLMs). Experiments with general-purpose LLMs in agricultural settings suggest that these models can provide useful assistance but often default to generic advice, lacking in local specificity Ibrahim et al. (2023); Calone et al. (2024). Agriculture-specific chat-bot systems such as Farmer.Chat Singh et al. (2024) that have multilingual support are promising, however there has been no quantitative evaluation of such systems. On the other hand, the handful of Agriculture-specific Question Answer (QA) datasets are only for high resource languages such as English, German and Chinese Dofitas et al. (2025); Jiang et al. (2025); Kästing and Hänig (2024); Kpodo et al. (2024).
In this paper, we present a novel synthetic QA dataset in three languages - English, Hindi and Punjabi, for the agriculture domain. In addition, we provide human-curated evaluation datasets for all three languages. We carry out extensive evaluations to determine the best LLM, as well as the prompt. In contrast to previous research that used either metrics such as BLUE and ROGUE Johnny and Jaya Nirmala (2022); Jiang et al. (2025) or LLM-as-judge Yang et al. (2024); Gauba et al. (2025) approaches for evaluation, we employ human evaluation at each step. Through this, we highlight the inadequacy of automated evaluation schemes for domain, as well as country-specific QA evaluation tasks. With the synthetic training data, we carry out two different experiments (i) train and test LLMs with language-specific data (ii) train with English data and use translate-test approachAhuja et al. (2023), where the questions in Hindi or Punjabi were first translated into English, fed to English trained model, and the generated responses were then translated back into Hindi or Punjabi, respectively. Our results demonstrate notable improvements from language-specific fine-tuning and further gains using the translate-test approach via an instruction-tuned English model. In Punjabi, factuality improved by 7.4% with fine-tuning and further increased by 8.7% using the translate-test approach. In Hindi, factuality rose by 10.8% through fine-tuning and gained an additional 10.1% with the translate-test method. The English model also demonstrated substantial improvements, with factuality increasing by 11.5%, and relevancy rising by 25.2% after fine-tuning.
These consistent improvements across all metrics underscore two key insights: (1) domain-specific fine-tuning significantly boosts performance, and (2) instruction-tuned English models, when used with translate test approach, outperform native language fine-tuned models
2 Dataset Creation
2.1 Synthetic Training Data Generation
2.1.1 Data Source
The agricultural data used to generate QA pairs was collected through web scraping from Vikaspedia [^1] https://en.vikaspedia.in/, an online encyclopedia created by the Government of India. A detailed summary of data source files and token statistics across these domains is presented in Table 1222Token counts computed using the LLaMA 3 8B Instruct tokenizer..
Topic Files Tokens
Agri Exports 9 24,813
Agri Inputs 60 105,348
Agri Insurance 9 36,643
Agro enterprises 31 101,939
Best Practices 75 174,997
Crop Production 637 2,250,709
Fisheries 112 544,672
Forestry 18 69,916
ICT applications_in_Agriculture 18 36,664
Livestock 107 415,639
Market information 13 32,483
National Schemes_for_Farmers 17 72,449
Policies and Schemes 86 213,880
Post Harvest Technologies 41 158,399
Poultry 25 75,997
State specific schemes for farmers 50 117,642
Women and agriculture 42 87,372
Total 1350 4,519,562
Table 1: Token and topic statistics across different topics in the Vikaspedia dataset.
The raw text in each file was split into manageable chunks of 2000 characters each, with an overlap of 200 characters between consecutive chunks. These chunk and overlap sizes were chosen based on experimentation to ensure that important contextual information was retained across chunks.
Instruction:
Suppose you are a knowledgeable expert in agriculture.
Please generate as many unique sets of corresponding question and answer pairs strictly based on the provided Context.
No additional context is required.
Context:
{context}
Response:
Question: [Question here]
Answer: [Answer here]
Examples:
{examples}
Figure 1: Prompt used for synthetic data generation. When prompting the model, the context tag will be replaced with a fixed input chunk from the agricultural dataset.
2.1.2 QA Generation Using LLMs
Paula et al. (2024) demonstrated that QA generation by prompting an LLM in a few-shot setting is more effective than approaches such as end-to-end generation (the sequence of questions and answers is generated given the input context) and pipeline methods (the answers are extracted first and questions are then generated in an answer-aware manner) and was therefore used for QA generation in our work.
DeepSeek-R1-Distill-Llama-70B (Guo et al., 2025) was used to create synthetic QA pairs, after preliminary experiments that involved Llama-3.1 405B as well (see Table 2 for sample QA pairs generated by each LLM).
As shown in Figure 1, the instruction used in our prompt was designed to generate QA pairs from a given text segment of fixed chunk size using few-shot prompting. This prompt design was inspired by several past research (Balaguer et al., 2024; Yehudai et al., 2024). Contrary to previous research that instructed the LLM to generate a fixed number of questions Paula et al. (2024), we instruct it to generate as many questions as possible. This caters to the varying information density in the text chunks.
A total of 60,140 question-answer pairs were created from the agricultural dataset in English. This data set was further translated into Hindi and Punjabi, using IndicTrans2 Gala et al. (2023).
Model Question and Answer
LLaMA Q: How can I manage thrips in groundnut?
A: Apply Fipronil 2 ml/litre.
DeepSeek Q: What should I do if my groundnut crop shows yellowish-green patches and brown necrotic areas?
A: This could be a sign of thrips infestation. Apply Fipronil 2 ml/litre to control the infestation.
Table 2: Comparison of a single QA pair generated by LLaMA and DeepSeek models.
2.2 Evaluation Dataset
We manually created an English test dataset by collecting QA pairs from multiple sources such as Tamil Nadu Agricultural University [^3] http://www.agritech.tnau.ac.in/expert_system/, Union Bank of India-FAQ-Agri-Products [^4] https://www.unionbankofindia.co.in/pdf/faq-agri-products.pdf, ICAR-NRRI, Cuttack [^5] https://icar-nrri.in/faq/, and ICAR-NRCP, Solapur [^6] https://nrcpomegranate.icar.gov.in/FAQ. From these sources, we manually generated 1,699 QA pairs, ensuring diversity in topics, relevance to agricultural advisory, and high-quality reference answers.
Additionally, a Hindi test dataset was developed by transcribing videos from the DD Kisan youTube channel [^7] https://www.youtube.com/watch?v=MYkBswpxRcA. 50 videos were selected at random to cover diverse agricultural topics. GenAI 2.0 Flash  Google DeepMind (2025) was employed to generate QA pairs from 500 token segments of the transcript, resulting in 1,200 Hindi QA pairs. Subsequently, these were translated into Punjabi using IndicTrans2 Gala et al. (2023). The Hindi and Punjabi data sets were manually cleaned and verified to ensure an accurate and high quality test data set.
3 Implementation Details
3.1 Experiments
Selection of LLMs: The LLMs selected for initial zero-shot experiments are shown in Table 3.
Punjabi Hindi English
Llama 3.1 Llama 3.1 Airavata
Llama 3 Llama 3 Llama 3
Navarasa Navarasa Deepseek
Mistral Airavata
Gemma
Table 3: Models selected for each language. The best performing models for each language were fine-tuned and are highlighted in bold.
Selection of Prompt: We started with the prompt provided by Silva et al. (2023) which was originally designed for agriculture examination contexts. However, our long-term goal is to support farmers with advisories. Hence, we adapted the prompt and generated six different prompt variations (see Appendix A) . The intuition behind each prompt is given in the Appendix B. Zero-shot experiments were carried out to select the best prompt which was used for further fine-tuning of models in all languages.
LLM Fine-tuning: The best performing model under the zero-shot setting was fine-tuned with the synthetic dataset we created. Further implementation details are provided in Appendix F, and the hyperparameters are listed in Appendix E. Fine-tuned models for each language were evaluated using their respective language-specific evaluation datasets. Further, we also conducted an experiment to evaluate the efficacy of the translate-test approach for Hindi and Punjabi languages.
3.2 Evaluation Metrics
The responses were evaluated by human evaluators using three main metrics: Agricultural Consensus, Relevancy, and Factuality. These metrics are inspired by similar evaluation methods used in prior research Singhal et al. (2023); Balaguer et al. (2024). More details about these metrics are provided in Table 4. Human evaluators followed evaluation guidelines as outlined in Appendix  C. Evaluators are native speakers of Hindi or Punjabi and are also fluent in English.
Multiple raters/measurements Koo et al. (2016) was used to calculate inter-annotator agreement scores for relevancy and factuality. Fleiss Kappa  Elangovan et al. (2025) was used for reliability. Justification for the selection of these measures and how to interpret the results are discussed in Appendix H.
No. Parameter Evaluation Question Scale
1 Agricultural Consensus (Con) How does the answer relate to the consensus of the agricultural domain? Yes / No
2 Relevancy (Rel) Does the answer provide information pertinent to the question, regardless of whether the facts are entirely accurate? 1 – 5
3 Factuality (Fact) Is the response generated correct, partially correct, or incorrect? Does it contain any content it shouldn’t? Correct / Incorrect / Partially Correct
Table 4: Evaluation criteria for human assessment
4 Results
Due to the scarcity of human evaluators, it was not possible to evaluate all the LLMs across all the data samples in the test set. Therefore, a subset of QA pairs was selected. From each QA pair, the question was used in the prompt, and the corresponding answer generated by the LLM was evaluated against the ground-truth answer by a human. Table 5 shows the amount of human evaluations carried out for the purpose of LLM selection.
4.1 Model and Prompt Selection
Setting Models/ Temp Unique Total
Prompts QA pairs Evals
Multilingual Model Comparison
English 4 LLMs 3 20 720*
Hindi 5 LLMs 1 100 1500*
Punjabi 3 LLMs 1 100 900*
Prompt Selection (1 LLM)
English 6 Prompts 1 25 150
Base vs. Finetuned Evaluation
English 2 (Base+FT-EN) 1 300 600
Hindi 2 (Base+FT-HI) 1 240 480
Punjabi 2 (Base+FT-PA) 1 240 480
Translate Test Approach (via FT-EN model)
Hindi 1 (FT-EN) 1 240 240
Punjabi 1 (FT-EN) 1 240 240
Total Evaluations 5,310
Table 5: Overview of all experiments. Three temperature values (0.2, 0.5, 0.7) were tested only for English model selection. The best temperature identified in that step was used for other experiments.
*Each QA pair in the multilingual model comparison experiment was independently evaluated by 3 annotators. All other evaluations were conducted by a single annotator.
77footnotetext: *Each QA pair in the multilingual model comparison experiment was independently evaluated by 3 annotators.
The results in Table 6 show that LLaMA-3-8B-Instruct, LLaMA-3.1-8B-Instruct and LLaMA-3.1-8B-Instruct emerged as the best performing models for English, Hindi and Punjabi, respectively, at temperature T = 0.2. The experiments were also conducted at T =[0.5, 0.7] and results are in Appendix I. Prompt 1 (see Appendix A)was used for these experiments. Inter-rater agreement was found to be $>$ 0.62 for all cases showing substantial agreement among annotators (see Table 7) .
Language Model Name Consensus Relevancy Factuality
English Airavata-7B-Instruct 100% 58.3% 30.0%
DeepSeek-8B-Instruct 100% 33.3% 28.3%
LLaMA-3-8B-Instruct 100% 98.3% 66.6%
Mistral-7B-Instruct 98.3% 93.3% 30.0%
Hindi Gemma-7B-Instruct 90% 26.3% 14.6%
LLaMA-3-8B-Instruct 84% 25.3% 13.3%
LLaMA-3.1-8B-Instruct 96% 74.6% 35.6%
Mistral-7B-Instruct 100.0% 37.3% 5.6%
Navarasa-7B-Instruct 91.0% 58.6% 21%
Punjabi LLaMA-3-8B-Instruct 87% 17.6% 12.3%
LLaMA-3.1-8B-Instruct 100% 43.3% 24.3%
Navarasa-7B-Instruct 95.6% 29.3% 15.6%
Table 6: Human Evaluation Results for Model Selection Across Languages at Temperature $T=0.2$
Among the six tested prompts, Prompt 1 achieved the highest scores across all evaluation metrics and was therefore selected as the final instruction format. The evaluation results are shown in Table 8.
Metric Calculation English Hindi Punjabi
Consensus Fleiss’ $\kappa$ 1.0 1.0 0.95
Relevancy ICC 0.79 0.62 0.77
Factuality ICC 0.82 0.93 0.75
Table 7: Inter-rater agreement for three languages across evaluation criteria.
4.2 Base vs. Fine-Tuned Model Results
For each language, the model that provided the best result in the zero-shot setup was used for fine-tuning experiments. Results are reported in Table 9. A notable improvement was observed for relevancy and factuality across all three languages after fine-tuning, despite a drop in performance for consensus for Hindi and Punjabi. The results in Table 9 indicated that the English fine-tuned model in translate-test approach outperformed the Hindi and Punjabi fine-tuned models.
Prompt Number Consensus Relevancy Factuality
Prompt 1 89.3% 90.6% 70.6%
Prompt 2 89.3% 80.0% 65.3%
Prompt 3 89.3% 77.3% 66.6%
Prompt 4 89.3% 54.6% 58.6%
Prompt 5 89.3% 74.6% 60.0%
Prompt 6 89.3% 80.0% 58.6%
Table 8: Human evaluation results for the selected English prompts.
Lang Model Consensus Relevancy Factuality
BASE FT TT BASE FT TT BASE FT TT
En LLaMA-3 96.0% 97.7% 59.0% 84.7% 46.6% 58.1%
Hi LLaMA-3.1 98.5% 96.7% 97.92 64.7% 67.4% 71.53 28.1% 38.9% 49.03
Pa LLaMA-3.1 97.4% 94.5% 98.75 41.0% 61.3% 73.47 21.4% 28.8% 37.50
Table 9: Comparison of Base, Fine-Tuned (FT) and Translate-test with En-FT model (TT) Performance Across Languages
5 LLM-as-a-Judge Evaluation
In the above experiments, 5,310 human evaluations were done. Although human evaluations are the gold standard for assessing model performance, they can be time consuming, expensive, and prone to differences in judgment between reviewers  Tan et al. (2025). To find more scalable alternatives, we employed the "LLM-as-a-Judge" approach for evaluation. We used three instruction-tuned models, namely Llama-4-Scout-17B-16E-Instruct, Llama-4-Maverick-17B-128E-Instruct, and DeepSeek-R1-Distill-Llama-70B [^8] We excluded larger models like Llama 4 Behemoth (288B) and DeepSeek-R1 (685B) due to high computational costs and resource constraints.. We generated prompts, each corresponding to the evaluation metrics ie. consensus, factuality and relevancy. See Appendix  A for prompts), which were adapted from G-eval  (Liu et al., 2023). The same QA pairs used for human evaluation were also evaluated using LLM as the judge, across all languages.
For the relevance and factuality metrics, the Spearman correlation  Spearman et al. (1904) was used to assess the agreement between the LLM’s judgments and those of human evaluators. For the agriculture consensus metric, which involves binary classification, refer to Appendix J. The mean rating from multiple evaluators was computed for each QA pair, and this average score was then correlated with the LLM’s corresponding rating. Llama-4-Scout-17B-16E-Instruct was identified as the best-performing model based on correlation scores.
Lang. Model Rel. (Spearman $\rho$) Fact. (Spearman $\rho$)
En DeepSeek 0.25 0.10
L4-Scout 0.27 0.32
L4-Maverick 0.21 0.18
Hi DeepSeek 0.54 0.47
L4-Scout 0.50 0.53
L4-Maverick 0.45 0.46
Pa DeepSeek 0.33 0.16
L4-Scout 0.39 0.34
L4-Maverick 0.28 0.38
Table 10: Correlation between LLM judges and human evaluators across languages and metrics. Relevancy and Factuality use Spearman’s $\rho$.
6 Conclusion
In this work, we introduced a multilingual fine-tuning dataset in English, Hindi, and Punjabi for Indian agricultural context, addressing the need for regionally and linguistically relevant datasets in this domain. We also proposed an evaluation dataset that can serve as a benchmark for assessing the performance of LLMs on agricultural queries. Our experiments demonstrate that fine-tuned models significantly outperform their base counterparts. Furthermore, for Hindi and Punjabi queries, a translate-test approach where the queries are translated into English and processed using the English fine-tuned model, yields better performance than directly fine-tuning models in those languages. These findings underscore the potential of cross-lingual transfer and resource-efficient strategies for building multilingual agricultural advisory systems. As the next step, we plan to validate our models and system through field testing with farmers.
7 Limitations
There are several limitations to our work. First, the data set created and the methodology used in this study include only textual data. However, in real-world scenarios, farmers may have image-based queries such as pictures of crops or pests/diseases. It is not addressed by our current approach.Our future work will focus on building a multimodal and multilingual dataset that can handle both textual and visual input. It will also incorporate a wider range of Indian languages.Second, human evaluation was done on a subset of the dataset, due to the resource-intensive nature of the task. In future, a more comprehensive human evaluation across the entire dataset is necessary to ensure model’s efficacy and deployment. Another limitation is that is that our experiments were conducted using relatively smaller models (7B to 8B parameter range) due to limited computational resources. In future, the effect of using larger models on performance can be explored. Our current work does not account for dynamic contextual factors such as weather conditions or soil health, which are often important in generating accurate and dynamic advisories. Integrating such contextual data through various APIs in an agentic framework can enable adaptive and real-time responses and it is an important direction for our future work. The human annotators volunteered their time to contribute to the project, and no financial compensation was provided.
Appendix A Tested Prompts
This section provides the six tested prompts used for evaluating Mistral-7B in the Prompt Selection experiment. Each prompt was designed to optimize the quality of the response for agricultural question-answer generation.
  • Prompt 1 (Selected Prompt)
    As an experienced agronomist proficient in farming techniques, crop management, and disease-resistant crop cultivation, you are tasked with answering questions based on your expertise. Questions may be provided in various languages such as English, Hindi, Marathi, etc.
    Provide only the answer relevant to the question. Do not include any other information. The answer should be in the same language as the question.
  • Prompt 2
    You are an expert in agricultural science with over 25 years of experience in farming, crop management, and sustainable agricultural practices. Your expertise spans across various domains, including soil health, irrigation techniques, pest control, and the integration of modern technology in agriculture. Your task is to answer questions related to the agriculture domain in a clear, concise, and informative manner. Ensure your responses are well-researched, accurate, and tailored to the specific context of the question. Questions may be provided in various languages such as English, Hindi and Marathi. Provide only an answer relevant to the question.
  • Prompt 3
    You are an expert in the field of agriculture with decades of experience in farming, crop management, soil science, and agricultural technology. Your knowledge spans traditional and modern farming practices, and you are fluent in multiple languages, including English, Hindi, Marathi, and others. Your expertise allows you to provide accurate, practical, and culturally relevant advice tailored to the specific needs of farmers, researchers, and agricultural enthusiasts.Your task is to provide detailed, accurate, and actionable answers exclusively related to the agriculture domain. Your responses must be in the same language as the question, whether it is in English, Hindi, Marathi, or any other language. Ensure your answers are clear, concise, and directly address the query while incorporating relevant agricultural principles, practices, and innovations.
  • Prompt 4
    You are a seasoned agricultural expert with over 25 years of experience in farming, crop management, and sustainable agricultural practices. Your deep understanding of soil health, pest control, irrigation systems, and modern farming technologies has made you a trusted advisor to farmers, policymakers, and agricultural organizations worldwide. Your task is to answer a specific question related to the agriculture domain. The question will be provided to you, and your response should be detailed, accurate, and tailored to the context of the query. Ensure your answer is backed by scientific evidence and a clear understanding of the latest advancements in agriculture.
  • Prompt 5
    You are an expert in agriculture and tasked with answering questions related to agriculture.
    Please answer the following question with detailed, accurate, and insightful information based on your knowledge in the agricultural domain.
  • Prompt 6
    You are an expert agricultural advisor with deep knowledge of farming, soil health, pest control, and modern techniques. Provide a clear and well-explained response to the following question, ensuring accuracy, practicality, and ease of understanding. If multiple solutions exist, explain the best approach considering efficiency, cost-effectiveness, and sustainability. Factor in seasonal changes, soil conditions, and climate impact when relevant.
Appendix B Prompt Intuition
The following summarizes the design intuition behind the six tested prompts used for agricultural QA generation:
  • Prompt 1 (Selected): This prompt was Designed to deliver concise and contextually relevant answers in the same language as the input, ensuring minimal drift from the question’s intent.
  • Prompt 2: This prompt was designed to simulate a highly knowledgeable agricultural expert, embedding deep domain-specific authority into the model’s persona. The aim was to guide the model to generate precise, well-informed, and context-aware answers like domain experts.
  • Prompt 3: This prompt was designed to extract detailed, accurate, and context-sensitive responses by positioning the model as a multilingual agricultural expert. Emphasis was placed on reasoning, cultural relevance, and linguistic alignment to ensure comprehensive, well-researched answers tailored to diverse user needs.
  • Prompt 4: This prompt was designed to position the model as an advisor trusted by farmers and policymakers alike, having scientific expertise with leadership, so that it can provide context-rich, evidence-based answers rooted in both practical farming knowledge and awareness of modern agricultural advancements.
  • Prompt 5: This prompt was kept minimal to evaluate the model’s inherent capabilities without heavy role conditioning, by providing only a brief instruction focused on expected output quality.
  • Prompt 6: The intent of this prompt was to encourage comparative reasoning and decision-making within itself. Instruct it to consider multiple solutions and select the best one based on accuracy and practicality. Ensuring responses are optimized for real-world agricultural conditions, including seasonal, soil, and climate factors.
Appendix C Human Evaluation Framework
This appendix details the evaluation framework used to assess the quality of LLM-generated answers in the agricultural domain. The evaluation is conducted by agriculture experts based on the following three criteria:
  • Agricultural Consensus: Does the response correlate with commonly accepted agricultural knowledge?
  • Relevancy: Does the response demonstrate evidence of correct recall of knowledge (e.g., correct facts)?
  • Factuality: Is the generated response factually accurate, partially correct, or incorrect?
1. Agricultural Consensus
Responses are labeled as Yes or No depending on whether they are related to the agricultural domain.
  • Yes: Indicates the answer is clearly associated with agriculture or farming practices, regardless of its correctness.
  • No: Indicates the answer is irrelevant to agriculture.
Examples:
  • Question: How to improve soil condition?
    Predicted Answer: There are several ways to improve soil condition such as use of organic matter, crop rotation, and maintaining pH of soil.
    Agricultural Consensus: Yes
  • Question: What are some recommended places to purchase apples from?
    Predicted Answer: The Apple Store is the best place. They are very knowledgeable, sell and service all their devices, sell AppleCare and you can be hands-on with the computer before you buy it.
    Agricultural Consensus: No
2. Relevancy
Responses are rated on a scale of 1–5 based on their relevance to the question. A score of 5 means highly relevant and accurate; a score of 1 means minimal understanding or relevance.
Examples:
  • Question: How to improve soil condition?
    Predicted Answer: There are several ways to improve soil condition such as use of organic matter, crop rotation, and maintaining pH of soil.
    Relevancy Score: 5
  • Question: What are some recommended places to purchase apples from?
    Predicted Answer: The State of Jammu and Kashmir and Himachal Pradesh are the leading apple producing States in India.
    Relevancy Score: 2
3. Factuality
Responses are compared with reference answers and categorized as:
  • Correct: Closely matches the reference answer.
  • Partially Correct: Contains some correct information but has inaccuracies.
  • Incorrect: Includes major inaccuracies or irrelevant information.
Examples:
  • Question: How to improve soil condition?
    Reference Answer: Techniques include use of organic matter, mulching, avoiding over-tillage, crop rotation, using cover crops, balancing soil pH, and using appropriate fertilizers.
    Predicted Answer: There are several ways to improve soil condition such as use of organic matter, crop rotation, and maintaining pH of soil.
    Factuality: Correct
  • Question: Which season is suitable for apple plantation in Himachal Pradesh?
    Reference Answer: February to early May
    Predicted Answer: The best time to plant apple trees in Himachal Pradesh is between January and March.
    Factuality: Partially Correct
  • Question: What is the full form of GMO?
    Reference Answer: Genetically Modified Organisms
    Predicted Answer: General Motors Operation
    Factuality: Incorrect
Appendix D Contributors to Evaluation Dataset
The evaluation dataset used in this study for English, Hindi and Punjabi languages was curated by Kashish Mahendra, Vidhi and Sudhir Rajput. Our approach follows the principles of participatory research, and data creators have been offered coauthorship to compensate for their contributions.
Appendix E Hyperparameters
Table 11: Fine-tuning hyperparameters for LLaMA-3-8B-Instruct using QLoRA.
Parameter Value
Base Model LLaMA-3-8B-Instruct
LoRA Rank ($r$) 16
LoRA Alpha 16
Target Modules q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj
Batch Size 2
Gradient Accumulation Steps 4
Max Sequence Length 2048
Learning Rate 2e-4
Epochs 1
Optimizer AdamW (8-bit)
Quantization 4-bit (bnb)
Checkpointing Enabled (Unsloth)
Scheduler Linear with warm-up
Tracking Weights & Biases
Tokenizer LLaMA 3 Instruct tokenizer
Table 12: Fine-tuning hyperparameters for LLaMA-3.1-8B-Instruct using QLoRA.
Parameter Value
Base Model LLaMA-3.1-8B-Instruct
LoRA Rank ($r$) 16
LoRA Alpha 16
Target Modules q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj
Batch Size 2
Gradient Accumulation Steps 4
Max Sequence Length 2048
Learning Rate 2e-4
Epochs 1
Optimizer AdamW (8-bit)
Quantization Not Used
Checkpointing Enabled (Unsloth)
Scheduler Linear with warm-up
Tracking Weights & Biases
Tokenizer LLaMA 3 Instruct tokenizer
Appendix F Implementation Details
We employed Parameter-Efficient Fine-Tuning (PEFT) using the QLoRA method Dettmers et al. (2023) for adapting both the LLaMA-3-8B-Instruct and LLaMA-3.1-8B-Instruct models. PEFT was implemented via the Unsloth framework, which enables memory-efficient training through 4-bit quantization (used only for LLaMA-3) and Low-Rank Adaptation (LoRA). Gradient checkpointing was enabled to reduce memory usage during backpropagation.
Fine-tuning followed an Instruct-style prompt format and was performed using Hugging Face’s SFTTrainer utility.
Appendix G Hardware and Infrastructure
All experiments were conducted on high-performance GPU nodes equipped with NVIDIA Tesla V100 (32GB) and NVIDIA A100 (80GB) GPUs. The environment was configured to efficiently support large-scale language model inference and fine-tuning, leveraging mixed precision and distributed processing where applicable.
Appendix H Inter-annotator Agreement Calculation
Since the same fixed set of evaluators assessed a specific, non-random sample set for each language, the two-way mixed effects, absolute agreement, multiple raters/measurements model was used to calculate agreement scores  Koo et al. (2016). According to standard interpretation, ICC values below 0.5 indicate poor reliability, values between 0.5 and 0.75 suggest moderate reliability, values between 0.75 and 0.9 denote good reliability, and values above 0.9 reflect excellent reliability. According to the interpretation scale proposed by Landis and Koch (1977) , the strength of agreement for Kappa statistic can be categorized with values less than 0 indicate Poor agreement, 0–0.20 as Slight, 0.21 – 0.40 as Fair, 0.41 – 0.60 as Moderate, 0.61 – 0.80 as Substantial, and 0.81 – 1.00 as Almost Perfect. For binary judgments related to agricultural consensus, Fleiss Kappa  Elangovan et al. (2025) was used to evaluate the reliability between evaluators.
$\text{ICC}=\frac{MS_{R}-MS_{E}}{MS_{R}+\frac{MS_{C}-MS_{E}}{n}}$ (1)
$MS_{R}$ = row mean square (subjects), $MS_{C}$ = column mean square (raters), $MS_{E}$ = error mean square, $n$ = number of raters per subject.
Appendix I Model Temperature Selection
Model T = 0.2 T = 0.5 T = 0.7
Con Rel Fact Con Rel Fact Con Rel Fact
DeepSeek R1 100.0% 33.3% 28.3% 100.0% 80.0% 31.6% 100.0% 70.0% 21.6%
LLaMA 3 100.0% 98.3% 66.6% 100.0% 93.3% 56.6% 100.0% 95.0% 51.6%
Mistral 100.0% 93.3% 35.0% 100.0% 75.0% 45.0% 57.0% 66.6%  55.0%
Airavata 100.0% 58.3% 30.0% 100.0% 95.0% 41.6% 100.0% 90.0% 40.0%
Table 13: English Model evaluation across 3 different temperature ($T$) settings using metrics: Agriculture Consensus (Con), Factuality (Fact), and Relevancy (Rel). The highest percentage is in bold.
Appendix J LLM-as-a-Judge : Pearson Correlation for Agriculture Consensus Metric
For the agriculture consensus metric, which involves binary classification, Pearson correlation Pearson et al. (1895); Fu et al. (2023) was considered. However, the correlation yielded a NaN value due to the absence of variance in the responses: both the annotators and the LLM provided identical labels for all instances, indicating complete agreement.
Appendix K LLM as Judge Prompts
Prompt for Agriculture Consensus Instruction: You will be given a question, a reference (gold) answer, and an LLM-generated answer. Your task is to rate the LLM-generated responses on one metric. Please make sure you read and understand these instructions carefully. Evaluation Criteria: Agricultural Consensus (Yes,No) - To assess the results generated by the LLM, we categorize them with labels of Yes or No depending on their correlation with the agriculture sector. A Yes label denotes a clear association with agricultural or farming activities, regardless of the accuracy of the response. Conversely, a No label indicates a lack of relevance to agricultural domain knowledge. NOTE: The generated response may not be accurate, but it should be relevant to the agriculture field. Evaluation Steps: 1. Understand the Purpose: - Recognize that the goal is to determine whether the LLM-generated response is relevant to the agriculture or farming domain, not whether it is factually accurate. 2. Read the Question: - Review the question to understand the context in which the LLM was expected to respond. This can provide cues about whether the topic is related to agriculture. 3. Analyze the Reference Answer: - Briefly examine the reference (gold) answer to see what type of response is expected and whether it is centered around agricultural knowledge or practices. 4. Review the LLM-Generated Answer: - Carefully read the generated answer to determine the domain it is addressing. Identify key terms, phrases, or themes that indicate its association (or lack thereof) with agriculture or farming. 5. Assess Domain Relevance: - Check for direct or indirect connections to agricultural topics such as crops, soil, irrigation, farming techniques, pest control, agricultural policy, rural economy, livestock, climate effects on farming, etc. 6. Make a Binary Decision: - If the generated answer clearly pertains to the agriculture sector, assign the label "Yes". - If the generated answer is unrelated to agriculture, farming, or associated practices, assign the label "No". 7. Avoid Judging Accuracy: - Do not consider factual correctness when assigning the label. The only criterion is whether the content is related to the agriculture domain. Inputs: Question: {question} Reference (Gold) Answer: {answer} LLM-Generated Answer: {answer_llm}
Prompt for Factuality Evaluation Instruction: You will be given a question, a reference (gold) answer, and an LLM-generated answer. Your task is to rate the LLM-generated responses on one metric. Please make sure you read and understand these instructions carefully. Evaluation Criteria: Factuality (Correct, Partially Correct, Incorrect) — The response generated by the LLM will be assessed against the reference answer to evaluate its accuracy. It will be classified as either "Correct", "Partially Correct", or "Incorrect". The Correct designation denotes a response that closely matches the reference answer, while an Incorrect label indicates significant inaccuracies or inappropriate information. Partially Correct refers to a response that is not entirely accurate but includes some elements or pieces of information that are correct. Evaluation Steps: 1. Understand the Question: - Read the question carefully to ensure a proper understanding of what is being asked. Clarify any ambiguity in the question to accurately evaluate the generated response. 2. Review the Reference Answer (Gold Answer): - Thoroughly analyze the reference answer to understand the key details and factual information it presents. Identify the core points, facts, or context that must be preserved in the LLM-generated answer. 3. Review the LLM-Generated Answer: - Read the LLM-generated response closely. Check if the information presented aligns with the reference answer. Identify the factual elements in the response and compare them with the reference answer to spot any discrepancies or omissions. 4. Assess Factual Accuracy: - Compare each key detail in the LLM-generated answer with the reference answer. - If the LLM-generated answer fully matches the factual information provided in the reference answer, rate it as Correct. - If the LLM-generated answer has some elements that align with the reference answer but also contains inaccuracies or incomplete information, rate it as Partially Correct. - If the LLM-generated answer includes significant inaccuracies or irrelevant information, rate it as Incorrect. 5. Contextualize the Answer: - Ensure that the context of the answer (e.g., domain knowledge, specific terminology) is understood and correctly represented. If the response introduces concepts that are out of context or misleading, it should be rated as Incorrect. 6. Check for Detail and Completeness: - Evaluate the level of detail in the response. If critical facts or key elements are missing from the LLM’s response compared to the reference answer, it should be considered Partially Correct or Incorrect, depending on the extent of the missing information. 7. Final Rating: - Assign the final rating to the LLM-generated answer based on the accuracy of the information provided compared to the reference answer. Inputs: Question: {question} Reference (Gold) Answer: {answer} LLM-Generated Answer: {answer_llm}
Prompt for Relevance Evaluation Instruction: You will be given a question and an LLM-generated answer. Your task is to rate the LLM-generated responses on one metric. Please make sure you read and understand these instructions carefully. Evaluation Criteria: Relevance (1–5) — The response will be rated on a scale of 1 to 5 based on its relevance to the question. A rating of 5 indicates a highly relevant and well-aligned response, while a rating of 1 indicates that the answer is largely irrelevant or mismatched with the question. NOTE: The answer must align with the question asked, regardless of its factual accuracy. A response that incorporates all pertinent information and clearly addresses the core intent of the question will score higher. Evaluation Steps: 1. Understand the Question: - Read the question carefully to identify the core intent and expected scope of the answer. Note keywords and concepts that define what kind of information is being asked. 2. Read the LLM-Generated Answer: - Examine the generated answer in its entirety. Determine what topic it is addressing, what information it provides, and how closely it connects to the original question. 3. Match Answer to Question Intent: - Check whether the LLM-generated answer addresses the main point of the question. - Consider if the answer is directly responsive to the question, or if it strays into unrelated or tangential topics. 4. Evaluate Depth and Completeness: - Assess whether the response includes not just a relevant answer but also supporting details, context, or additional information that enhances its relevance. - Responses that include a broader context or deeper insight into the question’s topic may score higher. 5. Rate on a Scale from 1 to 5: - 5: Fully relevant and directly answers the question with complete, well-contextualized information. - 4: Mostly relevant and answers the question well, with minor gaps or slight vagueness. - 3: Moderately relevant; some parts address the question, but the answer is incomplete or only partially focused. - 2: Minimally relevant; contains limited information related to the question or is largely off-topic. - 1: Irrelevant or barely addresses the question; shows poor understanding or misunderstanding of the question asked. 6. Ignore Factual Accuracy: - Do not factor in correctness of facts. Focus strictly on alignment with the question’s subject and intent. Inputs: Question: {question} LLM-Generated Answer: {answer_llm}